OpenStudy (anonymous):

please i need your help Application of counting Two cards are drawn at random from an ordinary deck of 52 cards (no jokers). Find the probability that the 2-card hand contains the given card. b.No less than 1 diamond. @kropot72

OpenStudy (kropot72):

The probablity of 1 diamond in the two-card hand is: \[P(1\ \diamond)=\frac{13C1\times39C1}{52C2}=\frac{13\times39\times2}{52\times51}\] The probablity of 2 diamonds in the two-card hand is: \[P(2\ diamonds)=\frac{13C2}{52C2}=\frac{13\times12}{52\times51}\] The required probability is: P(1 diamond) + P(2 diamonds)= you can calculate